I'm not sure if this problem is in the right section, but my teacher says that this problem pertains to real numbers.

Find a value of "y" for which the expression "(1 - y)(3 - y)(6 - y) = 18" becomes true.

Many thanks in advance.
: I'm not sure if this problem is in the right section, but my teacher says that this problem pertains to real numbers.

Find a value of "y" for which the expression "(1 - y)(3 - y)(6 - y) = 18" becomes true.

Many thanks in advance.

Answer by nerdybill(1051)

y = 0
.
To check, plug it back in:
(1 - y)(3 - y)(6 - y) = 18
(1 - 0)(3 - 0)(6 - 0) = 18
(1)(3)(6) = 18
18 = 18